# Enhancing ECG Classification Robustness with Lightweight Unsupervised Anomaly Detection Filters

### v1

@@ III-C Experimental Setup
This experiment evaluated the performance of the UAD methods across two distinct tasks: Detection of ODD classes and of signals unsuitable for analysis due to noise. Using the PTB-XL dataset, four scenarios were created. In each scenario, one diagnostic superclass (MI, CD, STTC, or HYP) was designated as the OOD class. All samples containing the OOD class, including those where it co-occurred with in-distribution classes, were removed from the training data (folds 1-8). The validation (fold 9) and test (fold 10) sets included both in-distribution and OOD samples. Performance was measured using the Area Under the Receiver Operating Characteristic curve (AUC). Using the BUT QDB dataset, the data was split patient-wise to avoid data leakage. The training set included patients with only Class 1 and Class 2 noise labels (acceptable quality). The test set included patients exhibiting Class 3 noise (unsuitable for analysis). A 10-second window was labeled as unsuitable for analysis due to noise (anomalous) if at least 1 second within that window was annotated as Class 3.
@@
While Experiment 1 benchmarked the intrinsic detection performance of the UAD methods, it does not quantify the impact of such a filter on the overall diagnostic pipeline. The second experiment was therefore designed to assess the practical benefit of employing a UAD filter upstream of a standard multilabel classifier in a realistic deployment scenario involving both OOD classes and signals unsuitable for analysis due to noise.
A high-performing multilabel classification model was required for this evaluation. We selected the resnet1d_wang model, identified as the best-performing architecture on the PTB-XL diagnostic superclasses in the comprehensive benchmark by Strodthoff et al. [34]. This model achieves a macro AUC of 0.930, comparable to more complex state-of-the-art methods like L5G-Net (0.9357) and X-ECGNet (0.936). Their PyTorch implementation [35] was ported to TensorFlow. The classifier was trained in the same four scenarios as Experiment 1, using only the four in-distribution classes and excluding the OOD class from training, validation, and initial testing sets to assess in-distribution performance. Deep SVDD was selected as the filter mechanism, as it demonstrated the best average performance across both detection of OOD classes and of signals unsuitable for analysis due to noise in Experiment 1. The hyperparameter configuration yielding the top result for each specific scenario was utilized.
To simulate realistic conditions, the PTB-XL validation and test sets were modified to include calibrated amounts of noise using the nst script from PhysioNet [36]. The ’em’ noise record from the MIT-BIH Noise Stress Test Database was injected (see Figure 2). The Signal-to-Noise Ratio (SNR) used by nst is defined based on the peak-to-peak amplitude of QRS complexes. To ensure validity of the simulation, it was crucial to determine the appropriate SNR for injection by estimating the noise level present in real-world signals unsuitable for analysis (BUT QDB Class 3 samples). The estimation required a specialized wavelet denoising procedure designed to preserve R-peak morphology for the subsequent amplitude calculation. We employed the ’db6’ mother wavelet due to its morphological similarity to the QRS complex, good energy compaction and good time localization. Using a decomposition level of 11 (appropriate for the 1,000 Hz sampling rate), Donoho’s universal threshold [37, 38] was applied with soft thresholding. The noise standard deviation was estimated from the median absolute deviation of the first-level detail coefficient to ensure accurate noise variance estimation. This calibrated noise was applied to approximately 16.5% of the validation and test data, matching the prevalence of Class 3 noise windows (signals unsuitable for analysis) observed in our BUT QDB test set. Crucially, this injection ratio was applied across both the in-distribution and OOD samples within each scenario (HYP, CD, STTC, MI). This stratification ensures a sufficient representation of samples that are both unsuitable for analysis due to noise and belong to an unseen class, allowing for a robust evaluation of the system’s ability to handle these concurrent challenges.
A custom accuracy metric was defined to evaluate the combined system, accounting for both filtering and classification efficacy. An outcome was counted as correct only if:
  1. 1.
    An ECG sample that should be rejected (OOD or too noisy) was successfully rejected by the filter.
  2. 2.
    An ECG sample that should be accepted was accepted and correctly classified by the classifier.
All other outcomes were considered incorrect. The performance of the classifier-only approach and the integrated classifier-filter system were compared using this custom accuracy. The impact of the filter was further analyzed by plotting the custom accuracy against the rejection rate.
@@ IV Results
The first experiment involved a comprehensive NAS to optimize six UAD methods (Deep SVDD, MAD, AE, VAE, NF, and DDPM) across five distinct anomaly detection scenarios. Four scenarios involved identifying an unseen OOD diagnostic superclass (HYP, CD, STTC, or MI), and one scenario focused on detecting ECG signals unsuitable for analysis due to noise. The objective was to maximize detection performance (AUC) while maintaining computational efficiency, constrained by a 512k parameter limit.
@@
For the detection of signals unsuitable for analysis due to noise, all methods performed exceptionally well, with AUCs exceeding 0.95. Deep SVDD (0.992 $\pm$ 0.002) and AE (0.991 $\pm$ 0.001) yielded the best results. Given its superior average performance across both detection of OOD classes and of signals unsuitable for analysis due to noise, Deep SVDD was selected as the filter mechanism for the subsequent experiment.
The second experiment evaluated the practical utility of deploying a UAD filter (using the optimized Deep SVDD model) upstream of a standard multilabel classifier in a simulated deployment environment containing both OOD classes and injected realistic noise. Performance was evaluated using a custom accuracy metric, which balances the need to filter anomalies with the need to process valid data. This metric credits the system only if an anomalous sample is rejected or if an in-distribution sample is accepted and correctly classified.
Figure 4 illustrates the relationship between the system’s rejection rate and the resulting custom accuracy. The rejection rate (x-axis) is varied by adjusting the anomaly score threshold of the Deep SVDD filter. A stricter threshold increases the filter’s sensitivity, thereby increasing the rejection rate. The dashed horizontal lines indicate the baseline accuracy of the classifier-only system, corresponding to a 0% rejection rate. In all scenarios, as the rejection rate increases from zero, the custom accuracy rises sharply, significantly surpassing the baseline performance. This improvement is attributed to the successful filtering of OOD and noisy samples that the classifier would otherwise misclassify. However, the curves demonstrate a clear trade-off between robustness and utility. The accuracy peaks at an optimal rejection rate, typically between 20% and 40% across the different scenarios. In this optimal region, the benefit of correctly rejecting anomalies significantly outweighs the cost of occasionally rejecting in-distribution data. Beyond this optimum, the accuracy declines because the filter becomes overly aggressive, increasingly rejecting in-distribution samples that the classifier could have correctly processed, thus diminishing the overall utility of the system.
TABLE II compares the custom accuracy of the baseline classifier-only system with the integrated classifier-filter system operating at the optimal rejection rate. The baseline classifier exhibited relatively low custom accuracy, ranging from 0.460 (MI) to 0.509 (HYP). This highlights the significant negative impact of OOD and noisy samples on overall system reliability in a realistic setting. The introduction of the Deep SVDD filter substantially improved the custom accuracy across all scenarios. The largest improvement was observed in the CD scenario, where accuracy increased from 0.466 $\pm$ 0.006 to 0.676 $\pm$ 0.006, representing an improvement of 21.0 percentage points. Similar significant percentage point gains were observed for MI (+18.1), STTC (+15.6), and HYP (+15.4). These results demonstrate the efficacy of the proposed filtering approach in enhancing the robustness of ECG classification systems.
V Discussion
@@
TABLE II: Comparison of overall system robustness (custom accuracy) between the baseline (classifier only) and the integrated system (classifier + optimized Deep SVDD filter) in the realistic deployment simulation (Experiment 2). Results are reported as mean $\pm$ standard deviation. The integrated system performance is reported at the optimal rejection rate, i.e. the peak accuracy observed in Figure4. The superior approach in each scenario is highlighted in bold font.
UAD
Method
Custom Accuracy
HYP CD STTC MI
Classifier only
0.509
$\pm$ 0.007
0.466
$\pm$ 0.006
0.462
$\pm$ 0.002
0.460
$\pm$ 0.006
Classifier + Filter
0.663
$\pm$ 0.011
0.676
$\pm$ 0.006
0.618
$\pm$ 0.004
0.641
$\pm$ 0.007
This study systematically investigated the application of UAD as an independent filtering mechanism to enhance the robustness of ECG classification systems in resource-constrained environments. The objective was to identify inputs unsuitable for reliable analysis, specifically unseen CVD classes and signals unsuitable for analysis due to noise. Through a comprehensive benchmark utilizing NAS under strict computational constraints ($\leq$ 512k parameters), we demonstrated that an upstream UAD filter significantly improves the overall accuracy and reliability of a downstream classifier, achieving an improvement in custom accuracy of up to 21.0 percentage points.
V-A Interpretation of UAD Method Performance
The comprehensive benchmark (Experiment 1) revealed significant disparities in the efficacy of the investigated UAD methodologies. Deep SVDD demonstrated the most robust overall performance, achieving the highest AUC in the CD (0.769), STTC (0.678), and MI (0.692) scenarios. Its superiority likely stems from its geometric, one-class objective, mapping in-distribution data into a minimal-volume hypersphere, which bypasses weaknesses inherent in other methods.
Traditional reconstruction methods (AE and VAE) assume anomalous data is harder to reconstruct than normal data; however, recent work suggests this assumption is often violated [39], potentially explaining their lower performance in detecting subtle OOD CVDs. AE and VAE performed relatively well for Hypertrophy (HYP) (AUCs $\approx$ 0.78). HYP is often characterized by significantly increased QRS amplitudes [40], which likely fall far outside the training distribution—a scenario AEs are generally effective at detecting.
MAD demonstrated strong performance, particularly for HYP (AUC 0.811). By reframing reconstruction as a contextual infilling task, MAD is sensitive to contextual anomalies involving violations of local temporal dependencies. This aligns well with detecting secondary repolarization abnormalities (”strain patterns”) sometimes seen in HYP [40]. Although theoretically suited for STTC (also a contextual anomaly), MAD’s performance was moderate (AUC 0.645). This may be due to its potential weakness in detecting widespread anomalies; if the bidirectional context is also anomalous, MAD might successfully predict a masked point based on its correlated, anomalous neighbors, resulting in a missed detection.
NFs, performing explicit density estimation, also excelled in the HYP scenario (AUC 0.811), likely detecting the distributional shift caused by altered voltage characteristics. However, their performance for MI and CD was modest. NFs are susceptible to assigning high likelihoods to OOD data [41]. Furthermore, in high-dimensional data, likelihood calculation can be dominated by local correlations [42, 43], causing the model to deem a sequence probable based on plausible local wave shapes while failing to recognize anomalous global structures.
DDPMs exhibited the weakest performance across OOD CVD tasks. We utilized standard Gaussian noise, following Ho et al. [26]. However, since ECG power is concentrated in lower frequencies (0-50Hz) [44], Gaussian noise (with uniform spectral density) may insufficiently corrupt these low-frequency components [27]. Consequently, low-frequency ECG anomalies might survive the partial noising process and be preserved during denoising, resulting in a low reconstruction error and poor detection performance, even though high frequencies (>>70Hz) also carry clinical value [44].
In contrast to the varied performance in OOD CVD detection, all methods demonstrated excellent efficacy (AUC $\geq$ 0.957) in detecting ECG signals unsuitable for analysis due to noise. Severe noise artifacts manifest as gross, high-amplitude distortions that represent global point anomalies far outside the distribution of acceptable quality ECGs, readily detected by all evaluated UAD paradigms.
V-B Variability in OOD Detection Difficulty
The results in TABLE I reveal significant variability in detecting different diagnostic superclasses as OOD. HYP was the most tractable (best AUC $\approx$ 0.81), followed by CD (best AUC $\approx$ 0.77). Detecting MI and STTC proved substantially more challenging (best AUCs $\approx$ 0.69 and 0.68). This disparity is attributable to the distinctiveness of their ECG manifestations and the effects of clinical confounding.
The relative ease of detecting HYP and CD likely stems from their pronounced impact on the QRS complex (ventricular depolarization). HYP typically manifests as augmented QRS amplitudes [45, 46, 47]. CD, particularly bundle branch blocks (BBBs), involves significant prolongation of the QRS duration and fundamental alterations in QRS morphology [48]. These substantial changes in amplitude, duration, and wave shape create features qualitatively distinct from other superclasses, facilitating clearer separation in the feature space.
In contrast, STTC and MI primarily involve abnormalities in the ventricular repolarization phase (ST segment and T-wave) [49]. These changes are often more subtle and variable than depolarization changes. Furthermore, there is inherent clinical overlap, as MI manifests primarily through STTC features (e.g., ST elevation/depression) [49].
@@ V-B Variability in OOD Detection Difficulty
However, the primary driver of the difficulty in isolating STTC and MI as OOD is likely clinical confounding, where in-distribution pathologies mimic the OOD class. Repolarization abnormalities are not exclusive to STTC or MI; they frequently occur as secondary effects of HYP and CD. HYP can induce ”strain” patterns (downsloping ST depression and T-wave inversion) [40] that overlap with changes seen in STTC and MI. Likewise, conditions in CD, such as LBBB, can produce significant secondary ST-T changes [48]. In the context of UAD, this confounding is critical. When STTC or MI is held out, the training data still contains samples with HYP and CD. Consequently, the UAD model learns that the morphological features associated with HYP strain and secondary CD changes are part of the in-distribution manifold. When subsequently presented with true ischemic patterns (STTC/MI), the model struggles to identify them as anomalous because similar morphologies are already represented within its learned boundary of normality, directly accounting for the lower AUC scores observed.
V-C Clinical Implications and System Robustness
The integration of the Deep SVDD filter upstream of the 1D ResNet classifier demonstrated substantial improvements in system robustness (Experiment 2). In the presence of both OOD classes and calibrated real-world noise, the baseline classifier exhibited poor custom accuracy (0.460 to 0.509). This highlights the vulnerability of standard diagnostic models when deployed outside controlled environments, where they are prone to making erroneous, often high-confidence, predictions on inputs they were not trained to handle. The UAD filter significantly mitigated this vulnerability, yielding percentage points improvements in custom accuracy ranging from 15.4 (HYP) to 21.0 (CD).
This enhancement is critical for the safe deployment of automated ECG analysis, particularly on wearable devices where continuous monitoring increases the probability of encountering novel pathologies or severe noise artifacts. By rejecting inputs that cannot be reliably processed, the system inherently adheres to the precautionary principle fundamental to clinical practice. This mechanism significantly reduces the risk of compromised patient safety by mitigating the likelihood that erroneous, high-confidence predictions on OOD data could precipitate inappropriate or potentially harmful clinical decisions.
The analysis of the rejection rate versus custom accuracy (Figure 4) reveals a crucial trade-off. While increasing the rejection rate initially improves accuracy by filtering out anomalies, an overly aggressive threshold leads to the rejection of in-distribution samples, diminishing the system’s utility. The optimal operating point, observed between 20% and 40% rejection in this study, must be carefully calibrated based on the specific clinical application.
The magnitude of the observed improvement is naturally dependent on the prevalence of OOD and noisy samples in the test set. However, our experimental design aimed to provide a realistic estimation. We designated only one diagnostic superclass as OOD and injected noise into 16.5% of the samples, mirroring the prevalence of signals unsuitable for analysis due to noise (Class 3) observed in the BUT QDB ambulatory dataset. Furthermore, many commercially available wearable devices like smartwatches are trained to detect a narrower set of conditions (e.g., Atrial Fibrillation [50]). In such applications, the spectrum of potential OOD CVDs is considerably larger, suggesting that the positive impact of an upstream UAD filter might be even more pronounced than demonstrated in this study.
V-D Comparison with Related Work
This work distinguishes itself from prior research that primarily focuses on Uncertainty Quantification (UQ) [5, 11] or post-hoc analysis of supervised classifiers [7, 8] for OOD detection. While those methods are valuable, they inherently rely on the features learned by the supervised model. In contrast, our approach employs independent UAD models trained solely on in-distribution data, offering a modular filtering mechanism applicable to any downstream classifier. Furthermore, this study addresses the simultaneous detection of both unseen pathologies and ECG signals unsuitable for analysis due to noise under strict efficiency constraints, a combination often overlooked in the literature. The utilization of NAS to systematically optimize UAD architectures specifically for the performance-efficiency trade-off is crucial for deployment in resource-constrained ECG analysis.
V-E Limitations and Future Work
This study has several limitations. First, the evaluation of computational efficiency relied on parameter count as a proxy for resource usage. Practical deployment on wearable hardware requires direct measurement of power consumption, latency, and memory footprint. Future work will involve implementing the Pareto-optimal models on representative microcontroller units (MCUs) to quantify these metrics and explore optimization techniques such as quantization and pruning.
Second, the methodology for simulating realistic noise in Experiment 2 relies on estimating the SNR of BUT QDB Class 3 samples. Determining the true SNR of severely corrupted ambulatory signals is inherently challenging, introducing the possibility of miscalibration in the simulation. However, the impact on the study’s conclusions is likely minimal, given the very high efficacy (AUC $\geq$ 0.957) of the UAD methods in detecting these severe noise artifacts (TABLE I).
Third, the resampling of ECG windows to 512 timesteps (51.2 Hz) was an empirically driven trade-off necessary to manage the substantial computational demands of the extensive NAS process. Although diagnostic information contained in higher frequencies may have been attenuated, our pre-screening tests confirmed that this resolution led to no loss in performance compared to the original frequencies.
Fourth, the study utilized ECG as the sole modality. Incorporating contextual sensor modalities, such as accelerometer data, could improve the ability to distinguish between signals unsuitable for analysis due to noise and those representing unseen CVD classes. This differentiation would enable differential system responses (e.g., discarding a sample unsuitable for anaylsis due to noise vs. alerting a clinician about an unseen CVD).
Finally, opportunities exist to further enhance UAD performance. Future research could explore advanced techniques, such as training NFs on high-level semantic representations rather than raw input data to mitigate issues with local correlations [43], utilizing simplex noise instead of Gaussian noise for DDPMs to better corrupt low-frequency components [27], or combining complementary UAD methods within an ensemble framework.
VI Conclusion
In this paper, we addressed the critical challenge of ensuring the reliability of machine learning models for ECG analysis when deployed in real-world settings, where encountering unseen cardiovascular diseases and signals unsuitable for analysis due to noise is inevitable. We conducted a comprehensive benchmark and comparative analysis of six UAD methods, optimized via NAS under strict resource constraints suitable for wearable devices.
Our evaluation demonstrated that Deep SVDD offers the superior balance of efficiency and detection performance for identifying both OOD pathologies and severe noise artifacts, outperforming traditional reconstruction-based methods, MAD, NFs, and DDPMs in this application.
By integrating the optimized Deep SVDD model as an upstream filter to a standard diagnostic classifier, we achieved a significant enhancement in overall system robustness. In a realistic deployment simulation incorporating both OOD classes and calibrated ambulatory noise, the integrated system yielded improvements in accuracy of up to 21.0 percentage points compared to the baseline classifier-only approach.
This work underscores the critical role of independent UAD mechanisms in safeguarding the performance of automated ECG analysis systems, paving the way for safer and more clinically reliable, low-resource wearable ECG monitoring systems at scale. Future research will focus on hardware implementation, direct power consumption analysis, and the exploration of advanced techniques to incorporate context information and further improve detection capabilities.

### v2

@@ III-C Experimental Setup
The first experiment evaluated the performance of the UAD methods across two distinct tasks: Detection of ODD classes and of signals unsuitable for analysis due to noise. Using the PTB-XL dataset, four scenarios were created. In each scenario, one diagnostic superclass (MI, CD, STTC, or HYP) was designated as the OOD class. All samples containing the OOD class, including those where it co-occurred with in-distribution classes, were removed from the training data (folds 1-8) but not from the validation (fold 9) and test (fold 10) sets. Performance was measured using the Area Under the Receiver Operating Characteristic curve (AUC) to get a threshold independent assessment. Using the BUT QDB dataset, the data was split patient-wise to avoid data leakage. The training set included patients with only Class 1 and Class 2 noise labels. The test set included patients exhibiting Class 3 noise. A 10-second window was labeled as unsuitable for analysis due to noise (anomalous) if at least 1 second within that window was annotated as Class 3.
@@
The second experiment was designed to assess the practical benefit of employing a UAD filter upstream of a standard multilabel classifier in a realistic deployment scenario involving both OOD classes and signals unsuitable for analysis due to noise. A high-performing multilabel classification model was required. We selected the resnet1d_wang model, identified as the best-performing architecture on the PTB-XL diagnostic superclasses in the comprehensive benchmark by Strodthoff et al. [22]. The classifier was trained in the same four scenarios as Experiment 1, using only the four in-distribution classes and excluding the OOD class from training to assess in-distribution performance. Deep SVDD was selected as the filter mechanism, as it demonstrated the best average performance across both detection of OOD classes and of signals unsuitable for analysis due to noise in Experiment 1. The hyperparameter configuration yielding the top result for each specific scenario was utilized. To simulate realistic conditions, the PTB-XL validation and test sets were modified to include calibrated amounts of noise using the nst script from PhysioNet [23]. The ’em’ noise record from the MIT-BIH Noise Stress Test Database was injected. The Signal-to-Noise Ratio (SNR) used by nst is defined based on the peak-to-peak amplitude of QRS complexes. We determined the appropriate SNR for injection by estimating the noise level present in real-world signals unsuitable for analysis (BUT QDB Class 3 samples). For this, we used wavelet denoising with Donoho’s universal threshold and soft thresholding [24].This calibrated noise was applied to approximately 16.5% of the validation and test data, matching the prevalence of Class 3 noise windows (signals unsuitable for analysis) observed in our BUT QDB test set. Crucially, this injection ratio was applied across both the in-distribution and OOD samples within each scenario (HYP, CD, STTC, MI) to ensure a sufficient representation of samples that are both unsuitable for analysis due to noise and belong to an unseen class. A custom accuracy metric was defined to evaluate the combined system, accounting for both filtering and classification efficacy. An outcome was correct if valid and classified correctly, or if invalid (OOD CVD or Noisy) and rejected by the filter. The performance of the classifier-only approach and the integrated classifier-filter system were compared using this custom accuracy. The impact of the filter was further analyzed by plotting the custom accuracy against the rejection rate.
@@ IV Results
While all evaluated methods successfully identified signals unsuitable for analysis due to noise (AUC $\geq$ 0.957), differentiating unseen pathologies revealed distinct architectural biases. AE and VAE models likely underperformed because they generalized too well, accurately reconstructing subtle anomalies rather than flagging them [25]. MAD showed promise for local violations (e.g., Hypertrophy strain patterns [26]) but likely struggled when widespread anomalies corrupted the bidirectional context required for accurate infilling. Normalizing Flows appeared susceptible to assigning high likelihoods to OOD data based on plausible local correlations, failing to capture global structural defects [27, 28]. Finally, DDPMs exhibited the weakest performance. We hypothesize that standard white Gaussian noise failed to sufficiently corrupt the dominant low-frequency components of the ECG, allowing anomalies to survive the denoising process [17].
@@
The results in TABLE I also reveal significant variability in detecting different diagnostic superclasses as OOD. Detecting MI and STTC proved substantially more challenging than HYP or CD. We hypothesize that this disparity is attributable mostly to the effects of clinical confounding where in-distribution pathologies mimic the OOD class. Abnormalities during the ventricular repolarization phase (ST segment and T wave) are not exclusive to STTC or MI; they frequently occur as secondary effects of HYP and CD. In the context of UAD, this confounding is critical. When STTC or MI is held out, the training data still contains samples with HYP and CD. Consequently, the UAD model learns that the morphological features associated with secondary effects of HYP and CD are part of the in-distribution manifold. When subsequently presented with STTC or MI, the model struggles to identify them as anomalous because similar morphologies are already represented within its learned boundary of normality, directly accounting for the lower AUC scores observed.
Given its superior average performance across both detection of OOD classes and of signals unsuitable for analysis due to noise, Deep SVDD was selected as the filter mechanism for the subsequent experiment. The best parameter selection in terms of AUC from the NAS for each OOD type was as follows: CD and HYP: layers: 3, filter_0: 12, kernel_0: 11, stride_0: 1, filter_1: 24, kernel_1: 9, stride_1: 1, filter_2: 128, kernel_2: 3, stride_2: 1, latent dim.: 128; STTC: layers: 3, filter_0: 128, kernel_0: 3, stride_0: 2, filter_1: 128, kernel_1: 3, stride_1: 1, filter_2: 32, kernel_2: 3, stride_2: 1, latent dim.: 128; MI: layers: 3, filter_0: 16, kernel_0: 9, stride_0: 1, filter_1: 16, kernel_1: 5, stride_1: 1, filter_2: 16, kernel_2: 5, stride_2: 1, latent dim.: 512.
Figure 3 illustrates the relationship between the system’s rejection rate and the resulting custom accuracy. The rejection rate (x-axis) is varied by adjusting the anomaly score threshold of the Deep SVDD filter. A stricter threshold increases the filter’s sensitivity, thereby increasing the rejection rate. The dashed horizontal lines indicate the baseline accuracy of the classifier-only system, corresponding to a 0% rejection rate. In all scenarios, as the rejection rate increases from zero, the custom accuracy rises sharply, significantly surpassing the baseline performance by up to 21.0 percentage points (CD scenario). This improvement is attributed to the successful filtering of OOD and noisy samples that the classifier would otherwise misclassify. However, the curves demonstrate a clear trade-off between robustness and utility. The accuracy peaks at an optimal rejection rate, typically between 20% and 40% across the different scenarios. In this optimal region, the benefit of correctly rejecting anomalies significantly outweighs the cost of occasionally rejecting in-distribution data. Beyond this optimum, the accuracy declines because the filter becomes overly aggressive, increasingly rejecting in-distribution samples that the classifier could have correctly processed, thus diminishing the overall utility of the system.
V Conclusion
This study demonstrated that deploying a lightweight UAD filter upstream of an ECG classifier significantly enhances robustness. By benchmarking six architectures under strict constraints ($\leq$512k parameters), we identified Deep SVDD as the Pareto-optimal solution. In a realistic deployment simulation, this filter rejected unreliable inputs—both noise and unseen pathologies—improving diagnostic accuracy by up to 21.0 percentage points. Future work will focus on implementing the models on physical microcontrollers to measure real-world latency and power consumption, exploring options to improve UAD performance by e.g. using high-level semantic representations for NFs or utilizing simplex noise fo DDPMs, and incorporating multimodal data (e.g. accelerometer data) to differentiate between motion artifacts and physiological anomalies and thus allow different system responses in those two cases.
Acknowledgment
We thank the Bundesministerium für Forschung, Technologie und Raumfahrt (BMFTR) for their support under the project 16KISK221 (”Holistische Entwicklung leistungsfähiger 6G-Vernetzung für verteilte medizintechnische Systeme (6G-Health)”).